OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of threei

OpenStudy (anonymous):

\[-3+i3\sqrt3 \]

OpenStudy (anonymous):

\[r = a^2 + b^2\] \[\theta =\tan^{-1}(b/a)\]

OpenStudy (anonymous):

\[r^2 = 3^2 + (3\sqrt3)^2 = 36\]

OpenStudy (anonymous):

sorry, I made a typo when I was writing the formula for r, it should be: \[r =\sqrt{a^2 + b^2}\]

OpenStudy (anonymous):

Anyway, we have that r = 6

OpenStudy (anonymous):

We also get that theta = -60

OpenStudy (anonymous):

This, however, is not correct, and highlights a flaw of the arctan function.

OpenStudy (anonymous):

If we just plug in the values we got, we get this: \[6(\cos(-60) + isin(-60)) = 3 - i3\sqrt3\]

OpenStudy (anonymous):

Which is actually our answer times -1

OpenStudy (anonymous):

In order to rectify this, we add 180 degrees to theta.

OpenStudy (anonymous):

In the end, we have: \[-3 + i3\sqrt3 = 6(\cos(120)+i \sin(120))\]

OpenStudy (anonymous):

How did you get that theta=-60?